Anova: Single Factor SUMMARY Groups Count CAR1 CAR2 CAR3 CAR4 SumAverage Variance 178.9 19.87778 9.309444 11119.4 10.85455 10.15873 11241.1 21.91818 13.06164 14.49 7.894333 9 10 144.9 ANOVA Source of Varid Ss MS P-value F crit Between ( 815.10793 271.7026 26.61437 2.39E-09 2.8587916 Within Gr 377.7282 37 10.20887 Total 1192.836 40
Conduct a test of hypothesis that the mean mileage of makes 2 and 3 do not differ. Use the method for single means when σ is not known with the Error MS serving as the pooled variance Reminders: Test statistic t- difference of means / standard error of difference of means. The standard error of the difference equals square root of the sum of variances of the two means. The variance of each mean is estimated by the error mean square/number of observations in that mean 13. What is the value of the t test statistic for testing the hypothesis that makes 2 and 3 do not differ in mileage'? Example 24
Answer #1

Sp^2 = (10.15873 + 13.06164)/2 = 11.610185

t = (Xbar 2 - Xbar3)/(sqrt( Sp^2 (1/n2 + 1/n3))

= ( 10.85455-21.91818 ) / sqrt(11.610185 * (1/11 + 1/11 ))

= - 7.6148